OpenStudy (anonymous):

completing the square for -3x^2-6x+9 and graph a x-int and y-int and vertex point

OpenStudy (anonymous):

-3(x^2 + 2x ) + 9 = 0 -3(x^2 + 2x + 1) + 9 = -3 -3(x + 1)^2 +12 = 0 Vertex (-1,12)

OpenStudy (anonymous):

how do i find two points to graph?

OpenStudy (anonymous):

Pick any point... let's say x = 2 so y = -3x^2 - 6x + 9 so y = -3(2)^2 - 6(2) + 9 y = -12 - 12 + 9 y = -15 so (2,-15) Let's say x = 0 y = -3(0)^2 - 6(0) + 9 y = 0 - 0 + 9 y = 9 so (0,9)

OpenStudy (anonymous):

thank you!!